Solve the system of equations by the substitution method. (If the system is dependent, let y = c and enter a general solution in terms of c
{6x + 5y= 7
x − 3y= 5
(x, y) =

Answer by Boreal(15235)

6x+5y=7
x=3y+5
6(3y+5)+5y=7
18y+30+5y=7
23y=-23
y= -1
x=-3+5=2
(2, -1) ANSWER
check in first equation
6(2)+5(-1)=7